# GenMol: A Drug Discovery Generalist with Discrete Diffusion

### v1

@@ B.4 Goal-directed Hit Generation Results of GenMol Variants
Table 8: Goal-directed hit generation results of GenMol variants. The results are the means and standard deviations of PMO AUC top-10 of 3 runs. The best results are highlighted in bold.
  Oracle GenMol GenMol-ZINC250k GenMol-TR GenMol-NR
albuterol_similarity 0.932 $\pm$ 0.007 0.912 $\pm$ 0.008 0.895 $\pm$ 0.033 0.872 $\pm$ 0.032
amlodipine_mpo 0.804 $\pm$ 0.006 0.797 $\pm$ 0.031 0.802 $\pm$ 0.016 0.769 $\pm$ 0.029
celecoxib_rediscovery 0.826 $\pm$ 0.018 0.826 $\pm$ 0.012 0.821 $\pm$ 0.010 0.859 $\pm$ 0.008
deco_hop 0.953 $\pm$ 0.016 0.954 $\pm$ 0.019 0.945 $\pm$ 0.006 0.917 $\pm$ 0.009
drd2 0.995 $\pm$ 0.000 0.995 $\pm$ 0.000 0.995 $\pm$ 0.000 0.995 $\pm$ 0.000
fexofenadine_mpo 0.894 $\pm$ 0.028 0.893 $\pm$ 0.028 0.886 $\pm$ 0.017 0.875 $\pm$ 0.019
gsk3b 0.986 $\pm$ 0.003 0.988 $\pm$ 0.003 0.985 $\pm$ 0.003 0.985 $\pm$ 0.003
isomers_c7h8n2o2 0.934 $\pm$ 0.002 0.938 $\pm$ 0.003 0.934 $\pm$ 0.003 0.897 $\pm$ 0.016
isomers_c9h10n2o2pf2cl 0.833 $\pm$ 0.014 0.838 $\pm$ 0.014 0.830 $\pm$ 0.016 0.816 $\pm$ 0.025
jnk3 0.856 $\pm$ 0.016 0.856 $\pm$ 0.019 0.848 $\pm$ 0.016 0.845 $\pm$ 0.035
median1 0.397 $\pm$ 0.000 0.398 $\pm$ 0.001 0.397 $\pm$ 0.000 0.397 $\pm$ 0.000
median2 0.355 $\pm$ 0.003 0.352 $\pm$ 0.008 0.350 $\pm$ 0.006 0.349 $\pm$ 0.004
mestranol_similarity 0.981 $\pm$ 0.003 0.981 $\pm$ 0.003 0.980 $\pm$ 0.002 0.970 $\pm$ 0.004
osimertinib_mpo 0.876 $\pm$ 0.008 0.880 $\pm$ 0.000 0.876 $\pm$ 0.008 0.876 $\pm$ 0.003
perindopril_mpo 0.703 $\pm$ 0.009 0.708 $\pm$ 0.004 0.703 $\pm$ 0.006 0.697 $\pm$ 0.014
qed 0.942 $\pm$ 0.000 0.942 $\pm$ 0.000 0.942 $\pm$ 0.000 0.927 $\pm$ 0.000
ranolazine_mpo 0.821 $\pm$ 0.011 0.821 $\pm$ 0.023 0.818 $\pm$ 0.016 0.809 $\pm$ 0.009
scaffold_hop 0.628 $\pm$ 0.008 0.632 $\pm$ 0.001 0.621 $\pm$ 0.012 0.617 $\pm$ 0.002
sitagliptin_mpo 0.573 $\pm$ 0.050 0.596 $\pm$ 0.003 0.560 $\pm$ 0.037 0.573 $\pm$ 0.006
thiothixene_rediscovery 0.687 $\pm$ 0.125 0.686 $\pm$ 0.103 0.686 $\pm$ 0.121 0.650 $\pm$ 0.073
troglitazone_rediscovery 0.867 $\pm$ 0.022 0.827 $\pm$ 0.062 0.853 $\pm$ 0.035 0.801 $\pm$ 0.062
valsartan_smarts 0.797 $\pm$ 0.033 0.797 $\pm$ 0.036 0.797 $\pm$ 0.036 0.739 $\pm$ 0.043
zaleplon_mpo 0.569 $\pm$ 0.005 0.597 $\pm$ 0.008 0.569 $\pm$ 0.014 0.406 $\pm$ 0.002
Sum 18.208 18.207 18.091 17.641
B.5 Goal-directed Lead Optimization Results with ZINC250k
We provide the goal-directed lead optimization results of GenMol-ZINC250k in Table 9. As shown in the table, GenMol-ZINC250k shows comparable performance to GenMol, both showing superior performance to the baselines in Table 4.
@@ B.5 Goal-directed Lead Optimization Results with ZINC250k
Table 9: Lead optimization results (kcal/mol) of GenMol variants. The results are the docking scores of the most optimized leads. Lower is better, and the best results are highlighted in bold.
@@
  Target protein Seed score $\delta=0.4$ $\delta=0.6$
6
GenMol GenMol-ZINC250k GenMol GenMol-ZINC250k
-7.3 -10.6 -10.5 -10.4 -10.6
parp1 -7.8 -11.0 -11.2 -10.5 -10.5
-8.2 -11.5 -10.1 -9.5 -8.8
-6.4 -7.5 -7.9 -7.0 -
fa7 -6.7 -8.4 -8.2 -7.6 -7.2
-8.5 - - - -
-4.5 -12.9 -12.9 -12.1 -12.0
5ht1b -7.6 -12.3 -12.0 -12.2 -11.8
-9.8 -11.8 -10.5 -11.0 -11.0
-9.3 -10.8 -10.7 - -
braf -9.4 -10.2 -11.5 -10.2 -10.0
-9.8 -10.6 -10.6 -10.2 -10.0
-7.7 -10.0 -10.0 -9.5 -9.7
jak2 -8.0 -9.9 -10.0 -9.4 -9.2
-8.6 -9.5 -9.5 - -

### v2

@@ E.10 Examples of Generated Molecules
Table 16: Ablation study on goal-directed hit generation. The results are the means and standard deviations of PMO AUC top-10 of 3 runs. The best results are highlighted in bold.
  Oracle Attaching (A) A + Token remasking A + GPT remasking A + Frag. remasking (F) A + F + MCG
albuterol_similarity 0.872 $\pm$ 0.032 0.895 $\pm$ 0.033 0.908 $\pm$ 0.039 0.932 $\pm$ 0.007 0.937 $\pm$ 0.010
amlodipine_mpo 0.769 $\pm$ 0.029 0.802 $\pm$ 0.016 0.780 $\pm$ 0.032 0.804 $\pm$ 0.006 0.810 $\pm$ 0.012
celecoxib_rediscovery 0.859 $\pm$ 0.008 0.821 $\pm$ 0.010 0.847 $\pm$ 0.006 0.826 $\pm$ 0.018 0.826 $\pm$ 0.018
deco_hop 0.917 $\pm$ 0.009 0.945 $\pm$ 0.006 0.955 $\pm$ 0.005 0.953 $\pm$ 0.016 0.960 $\pm$ 0.010
drd2 0.995 $\pm$ 0.000 0.995 $\pm$ 0.000 0.995 $\pm$ 0.000 0.995 $\pm$ 0.000 0.995 $\pm$ 0.000
fexofenadine_mpo 0.875 $\pm$ 0.019 0.886 $\pm$ 0.017 0.905 $\pm$ 0.012 0.894 $\pm$ 0.028 0.894 $\pm$ 0.028
gsk3b 0.985 $\pm$ 0.003 0.985 $\pm$ 0.003 0.986 $\pm$ 0.003 0.986 $\pm$ 0.003 0.986 $\pm$ 0.001
isomers_c7h8n2o2 0.897 $\pm$ 0.016 0.934 $\pm$ 0.003 0.915 $\pm$ 0.008 0.934 $\pm$ 0.002 0.942 $\pm$ 0.004
isomers_c9h10n2o2pf2cl 0.816 $\pm$ 0.025 0.830 $\pm$ 0.016 0.820 $\pm$ 0.018 0.833 $\pm$ 0.014 0.833 $\pm$ 0.014
jnk3 0.845 $\pm$ 0.035 0.848 $\pm$ 0.016 0.840 $\pm$ 0.021 0.856 $\pm$ 0.016 0.906 $\pm$ 0.023
median1 0.397 $\pm$ 0.000 0.397 $\pm$ 0.000 0.396 $\pm$ 0.000 0.397 $\pm$ 0.000 0.398 $\pm$ 0.000
median2 0.349 $\pm$ 0.004 0.350 $\pm$ 0.006 0.353 $\pm$ 0.004 0.355 $\pm$ 0.003 0.359 $\pm$ 0.004
mestranol_similarity 0.970 $\pm$ 0.004 0.980 $\pm$ 0.002 0.980 $\pm$ 0.002 0.981 $\pm$ 0.003 0.982 $\pm$ 0.000
osimertinib_mpo 0.876 $\pm$ 0.003 0.876 $\pm$ 0.008 0.873 $\pm$ 0.003 0.876 $\pm$ 0.008 0.876 $\pm$ 0.008
perindopril_mpo 0.697 $\pm$ 0.014 0.703 $\pm$ 0.006 0.703 $\pm$ 0.000 0.703 $\pm$ 0.009 0.718 $\pm$ 0.012
qed 0.927 $\pm$ 0.000 0.942 $\pm$ 0.000 0.941 $\pm$ 0.000 0.942 $\pm$ 0.000 0.942 $\pm$ 0.000
ranolazine_mpo 0.809 $\pm$ 0.009 0.818 $\pm$ 0.016 0.824 $\pm$ 0.012 0.821 $\pm$ 0.011 0.821 $\pm$ 0.011
scaffold_hop 0.617 $\pm$ 0.002 0.621 $\pm$ 0.012 0.626 $\pm$ 0.005 0.628 $\pm$ 0.008 0.628 $\pm$ 0.008
sitagliptin_mpo 0.573 $\pm$ 0.006 0.560 $\pm$ 0.037 0.566 $\pm$ 0.010 0.573 $\pm$ 0.050 0.584 $\pm$ 0.034
thiothixene_rediscovery 0.650 $\pm$ 0.073 0.686 $\pm$ 0.121 0.677 $\pm$ 0.122 0.687 $\pm$ 0.125 0.692 $\pm$ 0.123
troglitazone_rediscovery 0.801 $\pm$ 0.062 0.853 $\pm$ 0.035 0.832 $\pm$ 0.052 0.867 $\pm$ 0.022 0.867 $\pm$ 0.022
valsartan_smarts 0.739 $\pm$ 0.043 0.797 $\pm$ 0.036 0.764 $\pm$ 0.038 0.797 $\pm$ 0.033 0.822 $\pm$ 0.042
zaleplon_mpo 0.406 $\pm$ 0.002 0.569 $\pm$ 0.014 0.586 $\pm$ 0.011 0.569 $\pm$ 0.005 0.584 $\pm$ 0.011
Sum 17.641 18.091 18.074 18.208 18.362
@@
Table 17: Ablation study on goal-directed lead optimization (kcal/mol) with $\delta=0.4$. The results are the mean docking scores of the most optimized leads of 3 runs. Lower is better and the best results are highlighted in bold.
@@
  Target protein Seed score Attaching (A) A + Token remasking A + GPT remasking A + Frag. remasking (F) A + F + MCG
-7.3 -8.2 -8.7 -10.4 -10.6 -10.6
parp1 -7.8 -8.3 -8.2 -11.5 -11.0 -11.0
-8.2 - -10.9 -11.1 -10.9 -11.3
-6.4 -7.2 -7.8 -8.1 -8.4 -8.4
fa7 -6.7 -8.3 -8.0 -8.1 -8.4 -8.4
-8.5 - - - - -
-4.5 - -12.8 -12.1 -12.9 -12.9
5ht1b -7.6 -11.9 -11.8 -12.1 -12.3 -12.3
-9.8 - -11.3 -11.2 -11.8 -11.8
-9.3 -9.7 -10.7 -10.8 -10.8 -10.8
braf -9.4 - -10.8 -10.1 -10.2 -10.8
-9.8 - -10.5 -10.6 -10.6 -10.6
-7.7 -8.6 -8.7 -10.2 -10.0 -10.2
jak2 -8.0 -8.9 -9.0 -10.1 -9.9 -10.0
-8.6 - -9.2 -9.8 -9.5 -9.8
Figure 7: Examples of generated molecules on de novo generation.
Figure 8: Examples of generated molecules on fragment-constrained generation of Eliglustat.